### v2

# C-DiffSET: Leveraging Latent Diffusion for SAR-to-EO Image Translation with Confidence-Guided Reliable Object Generation

Abstract
Synthetic Aperture Radar (SAR) imagery provides robust environmental and temporal coverage (e.g., during clouds, seasons, day-night cycles), yet its noise and unique structural patterns pose interpretation challenges, especially for non-experts. SAR-to-EO (Electro-Optical) image translation (SET) has emerged to make SAR images more perceptually interpretable. However, traditional approaches trained from scratch on limited SAR-EO datasets are prone to overfitting. To address these challenges, we introduce Confidence Diffusion for SAR-to-EO Translation, called C-DiffSET, a framework leveraging pretrained Latent Diffusion Model (LDM) extensively trained on natural images, thus enabling effective adaptation to the EO domain. Remarkably, we find that the pretrained VAE encoder aligns SAR and EO images in the same latent space, even with varying noise levels in SAR inputs. To further improve pixel-wise fidelity for SET, we propose a confidence-guided diffusion (C-Diff) loss that mitigates artifacts from temporal discrepancies, such as appearing or disappearing objects, thereby enhancing structural accuracy. C-DiffSET achieves state-of-the-art (SOTA) results on multiple datasets, significantly outperforming the very recent image-to-image translation methods and SET methods with large margins.
{strip}[Uncaptioned image]
Figure 1: Qualitative results of SAR-to-EO image translation using our proposed C-DiffSET framework (Confidence Diffusion for SAR-to-EO Translation). Rows 1, 2, and 3 display translations on the QXS-SAROPT, SAR2Opt, and SpaceNet6 datasets, respectively. Our proposed C-DiffSET achieve superior structural accuracy and visual fidelity compared to very recent methods.
1 Introduction
Satellite imagery plays a critical role in various applications, including surveillance, transportation, agriculture, disaster assessment, and environmental monitoring [81, 60, 70, 4, 67]. A significant portion of these applications relies on Electro-Optical (EO) imagery, which provides multi-spectral data captured by EO sensors. However, a fundamental limitation of EO imagery is its susceptibility to weather and lighting conditions, reducing its usability in scenarios such as cloudy weather or nighttime. In contrast, Synthetic Aperture Radar (SAR) offers robust sensing capabilities under all weather conditions without the need for light, making it ideal for nighttime operations. Despite these advantages, SAR imagery presents several challenges. SAR data suffers from speckle noise due to the backscattering of electromagnetic waves [13, 8, 79, 57]. Additionally, SAR images are typically grayscale, lacking the rich spectral and color information in EO images, making them difficult to be interpreted, especially for non-experts. To enhance the usability of SAR data, SAR-to-EO image translation (SET) has emerged as a promising research area in remote sensing, facilitating scene interpretation for tasks such as cloud removal [52, 41, 12] and change detection [39, 40, 38].
Recent advancements [75, 50, 25, 31, 32, 54, 14, 28] have demonstrated the potential of generating EO-like outputs from SAR images, offering improved textures and enhanced color information. However, several challenges persist: (i) The domain gap between SAR and EO images makes SET an ill-posed problem, exacerbated by the inherent noise in SAR images; (ii) The misalignment between SAR and EO datasets is prone to occur due to their different sensor platforms, satellite positioning shifts, or acquisition conditions, as shown in Fig. 2; (iii) Temporal misalignments (discrepancies) occur due to different acquisition times, also resulting in the spatial misalignment with differences in object appearance and seasons. This makes it very complicated the SET problem (Fig. 2); (iv) Finally, the scarcity of paired SAR-EO datasets makes it challenging for existing methods to learn the SET effectively, often leading to overfitting or unstable results.
To address these limitations, we firstly propose a novel framework that leverages a pretrained Latent Diffusion Model (LDM) [47, 11] to improve SAR-to-EO image translation, which is denoted as Confidence Diffusion for SAR-to-EO Translation (C-DiffSET). The advantages of utilizing the pretrained LDM for SET tasks are as follows: (i) Since EO images share visual characteristics with natural images, we fine-tune the pretrained LDM—trained on large-scale natural image datasets—to transfer its representation power to the SET task. This addresses the issue of limited SAR-EO paired data, as our framework can leverage the pretrained knowledge from natural image distributions; (ii) The pretrained LDM operates in a down-sampled latent space rather than the pixel domain, making it more robust to local spatial misalignments caused by imperfect alignment processes; (iii) The variational auto-encoder (VAE) [29] from LDM can also be applicable for SAR images. As shown in Fig. 4, despite the presence of significant speckle noise in SAR data, the VAE encoder effectively embeds SAR images into the same latent space as EO images, leveraging its denoising capability. This enables the smooth transfer of SAR latents as conditioning information in the reverse diffusion process, ensuring the generation of EO outputs with accurate pixel-wise correspondence to the SAR inputs.
Figure 2: Examples of misalignments in paired SAR-EO datasets. Left: Local spatial misalignments caused by sensor differences or acquisition conditions. Right: Temporal misalignments (discrepancies) where objects (e.g., ships) appear or disappear between SAR and EO images due to their different acquisition times.
In addition to utilizing the pretrained LDM as a foundation model for our SET task, we introduce a confidence-guided diffusion (C-Diff) loss to handle temporal discrepancies. The temporal discrepancies, such as objects appearing in only one modality, can introduce artifacts and hallucinated content in the generated EO images. To mitigate this, the U-Net [48] in our framework predicts both each predicted noise and its corresponding confidence map, which quantifies the pixel-wise uncertainty in the predicted noise. The U-Net receives a channel-wise concatenated SAR and noisy EO features as input, allowing it to jointly model both contexts. This setup enables the confidence map to reflect the inherent uncertainty caused by temporal discrepancies, guiding the C-Diff loss to adaptively reduce penalties in regions where SAR-EO misalignments likely occur. As a result, the generated EO images achieve high pixel-wise fidelity while minimizing artifacts and hallucinations. Our key contributions are summarized as:
  • We propose C-DiffSET, the first framework to fine-tune a pretrained LDM for SET tasks, effectively leveraging their learned representations to overcome the scarcity of SAR-EO pairs and local misalignment challenges.
  • We introduce a novel C-Diff loss that can guide our C-DiffSET to reliably predict both EO outputs and confidence maps for accurate SET.
  • We validate our C-DiffSET through extensive experiments on datasets with varying resolutions and ground sample distances (GSD) [61], including QXS-SAROPT [19], SAR2Opt [80], and SpaceNet6 [56] datasets, demonstrating the superiority of our C-DiffSET that significantly outperforms the very recent image-to-image translation methods and SET methods with large margins.
2 Related Work
2.1 Image-to-Image Translation
Image-to-image translation has been widely studied in various fields, such as image colorization [76, 77], style transfer [84, 24, 55], and image inpainting [22, 43]. These models typically rely on carefully constructed paired datasets across domains for effective training. For instance, Pix2Pix [24] introduced conditional GANs with an $l_{1}$ loss for domain-specific translations, while ASAPNet [55] proposed a lightweight framework for converting segmentation maps into realistic images, demonstrating abstract-to-detailed generation. However, acquiring paired datasets remains challenging, and various methods have been proposed to work around unpaired data constraints [1, 35, 27]. CycleGAN [84] addressed this limitation using a cycle-consistency loss to facilitate training without paired data. StarGAN [7] extended this idea to support multiple domain mappings within a single model, though its deterministic one-to-one approach limits flexibility. MUNIT [20] overcame this by decomposing images into domain-invariant content and domain-specific style representations, enabling diverse mappings while maintaining content integrity. Further addressing non-bijective translation, StegoGAN [69] introduced steganography into GAN-based models, enhancing semantic consistency in cases of domain mismatch by reducing spurious features in generated images without additional supervision.
Diffusion-based approaches. Diffusion models [18, 47, 11, 44, 17, 49, 33], have gained significant attraction in image-to-image translation due to their iterative denoising process, effectively modeling complex data distributions. Palette [49] utilizes conditional Denoising Diffusion Probabilistic Models (DDPMs) [18] in the pixel domain, achieving high-quality translations. Pixel-space operations often incur high computational costs for high-resolution images, motivating recent advances that utilize latent space diffusion to achieve greater efficiency. BBDM [33] diffuses directly from input to target latent space, eliminating the need for separate noise conditioning but imposing strict pixel-wise alignment requirements between input and target images. DGDM [71] is built on BBDM with a deterministic approach, where a translator network generates target features prior to diffusion, adding computational overhead and dependency on the initial translator’s performance.
Figure 3: Overall framework of our Confidence Diffusion for SAR-to-EO Translation (C-DiffSET).
2.2 SAR-to-EO Image Translation (SET)
Deep learning-based image-to-image translation methods [24, 55, 22, 20, 1, 35, 82, 33] have been widely adopted to remote sensing applications, including SAR-to-EO image translation (SET). However, SAR imagery poses unique challenges due to inherent characteristics like speckle noise from backscattering effects [13, 8, 79, 57, 59, 74, 21], complicating the translation process. To address these challenges, most SET studies employ GAN-based approaches [63, 5, 66, 32, 30, 9, 75, 72]. Conditional GAN-based models are commonly used to leverage noisy SAR information as a conditioning input, while CycleGAN-based models align the SAR and EO domains by enforcing cycle-consistency constraints. MCGAN [25] utilized a CycleGAN-based model, incorporating a mask vector with class labels to represent different regions. DTCDN [34] employed NICE-GAN [6] for joint SET and change detection tasks, improving feature alignment between domains. PLFM [52] introduced a multimodal model that integrates spatial and temporal data with SAR guidance images to produce clear EO outputs. Additionally, CFCA-SET [31] proposed a coarse-to-fine SAR-to-EO model that incorporates near-infrared (NIR) images during training to refine EO generation.
Diffusion-based approaches. Recently, a limited number of diffusion-based studies have been explored for SET, primarily divided into DDPM-based [3, 2] and LDM-based [28, 54, 14] methods. DSE [54] utilize the BBDM [33] to generate EO predictions for downstream tasks like flood segmentation. Expanding on this, cBBDM [28] integrates SAR inputs as conditional information in BBDM to improve translation quality, while CM-Diffusion [14] further enhances BBDM by conditioning on color features, helping to preserve spectral consistency in EO generation.
Despite these advancements, both GAN-based and diffusion-based methods encounter limitations stemming from the scarcity of paired SAR-EO datasets. GAN-based models frequently face with convergence challenges and mode collapse under limited data, while diffusion-based approaches struggle with overfitting and slow convergence due to the need for extensive training from scratch. To overcome these constraints, our C-DiffSET leverages a pretrained LDM as foundational model. By capitalizing on the extensive generative priors of large-scale text-to-image models, the C-DiffSET effectively adapts to the SET task, achieving robust translation with a reduced overfitting risk.
2.3 Latent Diffusion Models (LDMs)
Latent Diffusion Models (LDMs) [47, 11] extend the foundational DDPMs [18], which apply iterative noise addition to pixel-space images in a forward process, transforming them toward following Gaussian distribution. DDPMs are trained to reverse this process, step-by-step denoising to recover high-quality images. While effective for complex data modeling, DDPMs are computationally demanding in high-resolution applications due to their pixel-space processing. In contrast, LDMs address this by shifting the diffusion process to a learned latent space, significantly reducing memory requirements and computational costs. Using a VAE to encode images into a compact latent space, LDMs perform denoising directly on these representations, retaining high-resolution details with improved efficiency.
LDMs as foundation models. Open-source [62] releases of foundation LDMs, such as Stable Diffusion (v1.5, v2.1, v3.0) [47, 11], SDXL [45], both of which were pretrained on extensive text-to-image (T2I) datasets (i.e., LAION-5B [51]), have fueled progress across generative domains. These models offer powerful priors for image synthesis, and have been widely adapted through fine-tuning, particularly for natural image applications [68, 73, 26, 15]. Leveraging their strong generative capacity, we adapt a pretrained LDM for the SET task.
3 Methods
3.1 Overview of Proposed C-DiffSET
We utilize a paired SAR-EO dataset =$\mathcal{I}=\{(\mathbf{X},\mathbf{Y})\}$, where $\mathbf{X}\in\mathbb{R}^{C_{\text{sar}}\times H\times W}$ represents SAR images of $H\times W$ sizes and $C_{\text{sar}}$ channels and $\mathbf{Y}\in\mathbb{R}^{C_{\text{eo}}\times H\times W}$ denotes EO images of $H\times W$ sizes and $C_{\text{eo}}$ channels. Our goal is to generate a predicted EO image $\widehat{\mathbf{Y}}$ corresponding to the given SAR input $\mathbf{X}$. Fig. 3 illustrates the proposed Confidence Diffusion for SAR-to-EO Translation (C-DiffSET) framework, which comprises three key components: (i) the embedding of SAR and EO images into the latent space, (ii) the forward diffusion process, and (iii) the reverse diffusion process with the confidence-guided diffusion (C-Diff) loss. As described in Sec. 3.2, SAR image $\mathbf{X}$ and EO image $\mathbf{Y}$ are passed through the VAE encoder $\mathcal{E}_{\text{vae}}$, generating the latent features $\mathbf{z}_{x}\in\mathbb{R}^{C\times h\times w}$ (SAR feature) and $\mathbf{z}_{y}\in\mathbb{R}^{C\times h\times w}$ (EO feature). In the forward diffusion process (Sec. 3.3), noise $\bm{\epsilon}$ is added to the EO feature $\mathbf{z}_{y}$ over timesteps $t$s. During the reverse diffusion process, the U-Net $\psi$ takes the noisy EO feature $\mathbf{z}_{y}^{t}$, the timestep $t$, and the SAR feature $\mathbf{z}_{x}$ as conditional inputs. The Denoising U-Net $\psi$ predicts both the noise $\hat{\bm{\epsilon}}_{t}$ and a confidence map $\hat{\mathbf{c}}_{t}$ that quantifies pixel-wise uncertainty of the prediction. Finally, in Sec. 3.4, we describe the inference stage, where the predicted EO image $\widehat{\mathbf{Y}}$ is generated by reversing the noise-added process.
3.2 SAR and EO Latent Space Generation
We utilize the pretrained VAE from LDM for mapping input images from the pixel space to the latent space. The VAE is frozen during our experiments and serves as both an encoder $\mathcal{E}_{\text{vae}}$ and a decoder $\mathcal{D}_{\text{vae}}$ for input images. Since the LDM has been pretrained on large-scale natural image datasets, the VAE is designed to accept 3-channel RGB images as input. Fig. 4 shows the results of applying the VAE encoder $\mathcal{E}_{\text{vae}}$ and decoder $\mathcal{D}_{\text{vae}}$ to SAR and EO images, validating its applicability for both modalities.
EO latent space. Since EO images $\mathbf{Y}$ are naturally represented as 3-channel RGB inputs, they are directly passed through the VAE without modification. In our experiments, we confirmed that the reconstruction error $\|\mathbf{Y}-\mathcal{D}_{\text{vae}}\left(\mathcal{E}_{\text{vae}}\left(\mathbf{Y}\right)\right)\|_{2}$ is minimal, ensuring that the VAE accurately preserves the content and structure of EO images. This indicates that the VAE’s output provides an upper bound on the achievable performance of our framework, serving as a stable baseline for EO image generation.
SAR latent space. SAR data $\mathbf{X}$, depending on the satellite sensors, is available as single-polarization SAR images of 1-channel (HH or VV component) or full-polarization ones of 4-channel (HH, HV, VH, and VV components). For 1-channel SAR images, we repeat the channel three times to match the VAE’s input requirements. For 4-channel SAR data, we construct a 3-channel input by using HH, the average of HV and VH, and VV components. After passing these inputs through the VAE, we observed that the reconstruction error $\|\mathbf{X}-\mathcal{D}_{\text{vae}}\left(\mathcal{E}_{\text{vae}}\left(\mathbf{X}\right)\right)\|_{2}$ remained low, and the outputs appeared visually pleasing, as shown in Fig. 4. Additionally, we examined the VAE’s behavior under varying levels of speckle noise added to SAR inputs. The results demonstrate that the VAE’s reconstruction process adapts spatially to the noise level, effectively denoising the inputs while preserving structural details. This suggests that the VAE can embed SAR images into the same latent space as EO images without additional training, allowing SAR features to serve as conditioning inputs during the diffusion process. This ensures that pixel-wise correspondence between SAR and EO latent spaces is maintained throughout the generation process.
Figure 4: Results of applying the VAE encoder and decoder from LDM to EO and SAR images. The first row shows input images, including EO and SAR images with different levels of speckle noise. The second row presents the corresponding VAE reconstructions, illustrating that both EO and SAR images are accurately reconstructed despite noise variations.
3.3 Training Strategy for the Diffusion Process
The diffusion process consists of two key stages: a forward process that progressively adds noise to the target EO feature, and a reverse process that Denoising U-Net $\psi$ predicts and removes the added noise.
Forward process. Following the DDPMs [18] framework, noise is added to the target EO feature $\mathbf{z}_{y}$ over a sequence of timesteps $t\sim\mathcal{U}(T)$, where $\mathcal{U}$ is a uniform distribution and $T$ is a total number of timesteps. Specifically, at each timestep $t$, a noisy version of the EO feature $\mathbf{z}_{y}^{t}$ is generated by sampling:
$\mathbf{z}_{y}^{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{y}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\mathcal{N}$ is a Gaussian distribution, $\bm{\epsilon}\in\mathbb{R}^{C\times h\times w}$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ [18] determines the noise magnitude at each timestep $t$.
[…]
$[\hat{\bm{\epsilon}}_{t}\mid\hat{\mathbf{c}}_{t}]=\psi\left([\mathbf{z}_{y}^{t}\mid\mathbf{z}_{x}],\;\mathbf{z}_{c},\;t\right),$ (2)
[…]
Confidence-guided diffusion (C-Diff) loss. Seitzer et al. [53] proposed the $\beta$-NLL loss to capture aleatoric uncertainty for regression, classification and generative tasks. Inspired by the $\beta$-NLL loss [64, 53], we firstly adopt it into the training stage of diffusion process to enhance pixel-wise fidelity and to manage temporal inconsistencies between SAR and EO images. We apply the confidence output of $\psi$ at each diffusion step to the $\beta$-NLL loss for stable SET, which is called C-Diff loss $\mathcal{L}_{\text{C-Diff}}$. That is, $\mathcal{L}_{\text{C-Diff}}$ uses a learned confidence map $\hat{\mathbf{c}}_{t}$ to adaptively weight the predicted noise $\hat{\bm{\epsilon}}_{t}$ pixel-wise. $\mathcal{L}_{\text{C-Diff}}$ is designed to prioritize high-confidence areas while deweighting uncertain regions, thus improving robustness in regions with temporal misalignments (e.g., dynamic objects in SAR or EO images). $\mathcal{L}_{\text{C-Diff}}$ optimizes $\psi$ by combining a weighted pixel-wise reconstruction loss and a regularization term for the confidence map $\hat{\mathbf{c}}_{t}$:
$\mathcal{L}_{\text{C-Diff}}=\left\lVert\left(\bm{\epsilon}-\hat{\bm{\epsilon}}_{t}\right)\odot{\hat{\mathbf{c}}_{t}}^{\beta}-\log{\hat{\mathbf{c}}_{t}}^{\beta}+\tau\right\rVert_{2},$ (3)
where $\tau$ is a margin term ensuring non-negativity of the loss, and $\beta=1$. $\hat{\mathbf{c}}_{t}$ effectively acts as an adaptive weighting factor, allowing the model to focus more on well-aligned regions and reduce penalties in uncertain areas. The $\log$ term serves as a regularizer, preventing $\hat{\mathbf{c}}_{t}$ from collapsing to zero. $\mathcal{L}_{\text{C-Diff}}$ allows C-DiffSET to produce EO outputs of high structural accuracy while mitigating artifacts and hallucinations that can arise in temporally inconsistent regions.
3.4 Inference Stage for EO Image Prediction
Given an unseen SAR image $\mathbf{X}$, we pass it through the VAE encoder $\mathcal{E}_{\text{vae}}$ to obtain its SAR latent code $\mathbf{z}_{x}=\mathcal{E}_{\text{vae}}(\mathbf{X})$, which serves as the conditioning information for the reverse diffusion process. For inference, the Denoising U-Net $\psi$ iteratively refines the noisy latent code $\hat{\mathbf{z}}_{y}^{T}$, to reconstruct the target EO latent code $\mathbf{z}_{y}^{0}=\mathbf{z}_{y}$ by denoising it back to the clean EO latent code through noise prediction $\hat{\bm{\epsilon}}_{t}$. Starting from pure noise, $\hat{\mathbf{z}}_{y}^{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\psi$ iteratively refines the EO latent code $\hat{\mathbf{z}}_{y}^{t}$ by predicting the noise $\hat{\bm{\epsilon}}_{t}$ to be removed at each timestep $t$ as:
$[\hat{\bm{\epsilon}}_{t}\mid\mathsf{Dummy}]=\psi\left([\hat{\mathbf{z}}_{y}^{t}\mid\mathbf{z}_{x}],\;\mathbf{z}_{c},\;t\right),$ (4)
where $\mathsf{Dummy}$ indicates dummy confidence values. The predicted noise $\hat{\bm{\epsilon}}_{t}$ is then used to compute $\hat{\mathbf{z}}_{y}^{t-1}$ in the reverse diffusion process, following the formulation in [18], which progressively denoises the EO latent code until it converges to the target EO latent code. The final EO latent code $\hat{\mathbf{z}}_{y}^{0}$ is passed through the VAE decoder $\mathcal{D}_{\text{vae}}$ to generate the reconstructed EO image $\widehat{\mathbf{Y}}=\mathcal{D}_{\text{vae}}(\hat{\mathbf{z}}_{y}^{0})$.
Figure 5: Qualitative SAR-to-EO image translation on the SpaceNet6 dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
4 Experiment
4.1 Datasets
We evaluate our C-DiffSET framework on three publicly available SAR-to-EO datasets: QXS-SAROPT [19], SAR2Opt [80], and SpaceNet6 [56]. These datasets vary in satellite platforms, GSD [61], and SAR polarization modes, allowing us to assess the robustness and generalizability of our approach across diverse real-world scenarios.
QXS-SAROPT [19]. The QXS-SAROPT dataset contains 20,000 SAR and EO image pairs captured by the Gaofen-3 satellite (SAR) and Google Earth (EO). The SAR images are acquired in a single-polarization mode. The EO images consist of RGB channels, covering various port cities. Each image patch measures 256$\times$256 pixels with an 1-m GSD, focusing on complex maritime environments.
SAR2Opt [80]. The SAR2Opt dataset provides 2,076 SAR and EO image pairs obtained from the TerraSAR-X satellite (SAR) and Google Earth (EO). The SAR images are captured in a single-polarization mode. The corresponding EO images contain RGB channels and cover diverse Asian cities. Each patch measures 600$\times$600 pixels with an 1-m GSD, making this dataset particularly useful for urban area analysis and infrastructure monitoring.
SpaceNet6 [56]. The SpaceNet6 dataset offers SAR and EO image pairs captured by Capella Space (SAR) and Maxar WorldView-2 (EO) satellites. The SAR images are acquired with full-polarization, enabling detailed analysis of surface structures. The EO images include RGB and NIR bands, though only the RGB subset is used in our experiments. This dataset contains 3,401 SAR-EO image pairs, with each patch size of 900$\times$900 pixels and a 0.5-m GSD, focusing on urban landscapes and building detection tasks.
Figure 6: Qualitative SAR-to-EO image translation results on the SAR2Opt dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
  Types Methods Publications SAR2Opt Dataset SpaceNet6 Dataset
[…]
GANs Pix2Pix [24] CVPR 2017 196.87 0.426 0.0006 0.216 15.422 124.55 0.256 0.0102 0.522 19.357
CycleGAN [84] ICCV 2017 139.72 0.425 0.0022 0.224 14.931 114.81 0.274 0.0097 0.493 17.798
CFCA-SET [31] TGRS 2023 152.27 0.430 0.0009 0.223 15.183 164.78 0.279 0.0097 0.498 18.297
StegoGAN [69] CVPR 2024 144.54 0.398 0.0034 0.237 15.624 75.12 0.244 0.0106 0.516 18.958
LDMs BBDM [33] CVPR 2023 94.72 0.473 0.0005 0.234 15.131 81.86 0.302 0.0019 0.217 17.678
DGDM [71] ECCV 2024 156.12 0.541 0.0004 0.273 15.568 238.37 0.438 0.0015 0.253 17.124
cBBDM [28] arXiv 2024 97.64 0.394 0.0022 0.285 16.591 72.77 0.243 0.0079 0.254 19.033
C-DiffSET (Ours) - 77.81 0.346 0.0035 0.286 16.613 37.44 0.142 0.0151 0.567 21.022
Table 1: Quantitative comparison of image-to-image translation methods and SET methods on SAR2Opt and SpaceNet6 datasets. Red indicate the best performance in each metric.
4.2 Experiment Details
All experiments were implemented using PyTorch [42] and conducted on a single NVIDIA A6000 GPU. Each model was fine-tuned for 50,000 iterations, with a 100-step warmup period. We employed the AdamW optimizer [36] with an initial learning rate of $3\times 10^{-5}$ and a weight decay of 0.01. A cosine-annealing scheduler [37] was used to progressively reduce the learning rate at each iteration. For the pretrained LDM, we utilized Stable Diffusion v2.1 [47, 62], and the text encoder was frozen as the CLIP-ViT-H/14 [46, 23] text encoder. For the latent features, we set the spatial resolution parameters as $h=H/8$, $w=W/8$, and the channel dimension as $C=4$. The training noise scheduler was based on DDPM [18] with a total of 1,000 steps. For inference, we employed a more efficient DDIM [58] noise scheduler with total 50 inference steps to accelerate the generation process. Since SAR features were incorporated as additional inputs to the U-Net $\psi$, the weights of the $\psi$’s first convolutional layer were initialized by repeating the original weights across the SAR channels. Additionally, the confidence map was initialized to 1 to ensure stable optimization, starting from the standard $\ell_{2}$ loss, and $\tau=\log 2\pi$ [53] was applied. For the QXS-SAROPT dataset, we used the provided original 256$\times$256-sized patches without modification, with a batch size of 64. For the SAR2Opt and SpaceNet6 datasets, we performed random cropping to 512$\times$512-sized patches and set the batch size to 16. Data augmentation techniques, including random horizontal and vertical flips and random rotations (multiples of 90 degrees), were applied during training to improve generalization. To ensure reproducibility, a random seed of 2,025 was fixed for all experiments. Data was split with 80% used for training and 20% for testing across all experiments. We evaluated the performance of our framework using Fréchet Inception Distance (FID) [16], Learned Perceptual Image Patch Similarity (LPIPS) [78], Spatial Correlation Coefficient (SCC) [83], Structural Similarity Index (SSIM) [65], and Peak Signal-to-Noise Ratio (PSNR).
  Types Methods QXS-SAROPT Dataset
[…]
GANs Pix2Pix [24] 196.89 0.454 0.0000 0.247 14.924
CycleGAN [84] 195.38 0.455 0.0001 0.251 14.977
CFCA-SET [31] 79.06 0.406 0.0006 0.273 15.094
StegoGAN [69] 85.60 0.391 0.0019 0.280 15.580
LDMs BBDM [33] 65.15 0.522 0.0004 0.238 13.946
DGDM [71] 147.23 0.634 0.0001 0.288 11.564
cBBDM [28] 69.47 0.420 0.0023 0.304 16.248
C-DiffSET (Ours) 18.15 0.293 0.0108 0.372 18.077
Table 2: Quantitative comparison of image-to-image translation methods and SET methods on the QXS-SAROPT dataset. Red values represent the best results.
  Pretrained LDM $\mathcal{L}_{\text{C-Diff}}$ SAR2Opt Dataset
[…]
98.98 0.385 0.0014 0.255 15.815
78.14 0.349 0.0029 0.280 16.460
77.81 0.346 0.0035 0.286 16.613
Table 3: Ablation study on the SAR2Opt dataset evaluating the impact of pretrained LDM and confidence-guided diffusion loss.
4.3 Experimental Results
For comparative analysis, we used official implementations for general image-to-image translation methods [24, 84, 69, 33, 71]. For SET-specific methods [31, 28] (marked by $\dagger$), where official codes are generally unavailable due to the specialized nature of the field, we re-implemented each method according to their technical descriptions.
Qualitative comparison. As illustrated in Fig. 1, Fig. 5, and Fig. 6, the GAN-based methods generally struggle with stability during training, leading to prominent artifacts in SET. The LDM-based methods, such as BBDM [33] and cBBDM [28], face challenges due to their reliance on direct diffusion from SAR input $\mathbf{X}$ to EO output $\mathbf{Y}$. This setup makes them particularly susceptible to local spatial misalignments, producing blurred and incoherent results. DGDM [71] seeks to improve the initial stage of diffusion through a deterministic approach, employing a lightweight translation network to create a translated EO latent from SAR latent. However, the simplistic nature of this translation network fails to capture the intricate EO features, resulting in suboptimal initial latents for the diffusion process. In contrast, our C-DiffSET leverages pretrained LDM as foundational model, effectively addressing alignment issues via a confidence-guided diffusion loss, leading to yield sharper and more structurally coherent EO images, and to outperform both GAN-based and diffusion-based baselines in visual fidelity. Notably, as shown in Fig. 6, slight discrepancies between our generated EO images and the target EO images can occur. These differences arise from our model’s emphasis on structural alignment with the input SAR image, prioritizing SAR-EO correspondence over exact visual matching with the target EO.
Quantitative evaluation. In the quantitative assessment (see Table 1 and Table 2), the GAN-based methods yield lower PSNR, SSIM, and SCC scores, indicating their propensity for artifacts and limited spatial coherence. The LDM-based methods, BBDM and cBBDM, also perform inadequately due to their direct diffusion setup from SAR to EO, which exacerbates pixel-wise misalignment issues and inflates LPIPS and FID scores. The DGDM incorporates a lightweight translator network to initialize the diffusion process, but its simplistic structure fails to encapsulate detailed EO features, resulting in lower PSNR and SSIM scores. In contrast, our C-DiffSET achieves superior performance across all metrics, enabled by pretrained LDM and the confidence-guided diffusion loss, which jointly tackles SAR-specific noise and alignment challenges. This adaptation allows the C-DiffSET to attain the highest PSNR, SSIM, and SCC values, along with the lowest LPIPS and FID scores, highlighting its enhanced fidelity and perceptual quality in SET.
4.4 Ablation Studies
Impact of pretrained LDM and $\mathcal{L}_{\text{C-Diff}}$. To assess the contribution of the pretrained LDM and the confidence-guided diffusion loss $\mathcal{L}_{\text{C-Diff}}$, we conducted an ablation study on the SAR2Opt dataset, as shown in Table 3. The pretrained LDM provides strong initialization for EO generation, improving the baseline PSNR and SSIM scores when compared to training from scratch. Furthermore, incorporating $\mathcal{L}_{\text{C-Diff}}$ somewhat enhances these metrics and lowers LPIPS and FID values, improving both perceptual fidelity and structural consistency. Fig. 7 shows confidence maps generated on the training dataset at timestep $t=T/2$ (midpoint of the denoising process), highlighting the areas of temporal discrepancy in SAR-EO pairs. This loss enables the U-Net $\psi$ to down-weight uncertain regions where objects are temporally misaligned across modalities, thereby reducing artifacts and ensuring coherent EO outputs.
4.5 Limitations
Scalability across diverse satellites. The VAE in C-DiffSET is designed for 3-channel inputs, matching typical RGB images. This limits its direct application to datasets with more channels (e.g., RGB+NIR or multispectral data). An extension to supporting such inputs would require further exploration and tuning for effective adaptation.
Figure 7: Confidence maps generated by C-DiffSET at timestep $t=T/2$ on SAR-EO paired datasets: QXS-SAROPT, SAR2Opt, and SpaceNet6. Each row illustrates the input SAR image $\mathbf{X}$, the target EO image $\mathbf{Y}$, and the corresponding confidence map $\hat{\mathbf{c}}_{t}$.
5 Conclusion
In this work, we propose C-DiffSET, a novel framework that addresses key challenges in SAR-to-EO image translation. To the best of our knowledge, this is the first method to fully leverage LDM for SET tasks, mitigating issues caused by the limited availability of paired datasets. We introduce a C-Diff loss to handle temporal discrepancy between SAR and EO acquisitions, ensuring pixel-wise fidelity by adaptively suppressing artifacts and hallucinations. C-DiffSET achieves SOTA performance across datasets with varying GSD, demonstrating its effectiveness in real-world scenarios. Our framework provides a scalable foundation for future SET research and can be extended to other remote sensing applications and modalities.
Figure 8: Visualization of C-DiffSET results across inference timesteps on the SpaceNet6 dataset with a total of 50 inference steps. Each column corresponds to an inference step, starting with substantial noise at $t=981$ and progressively refining the output to match the target EO image at $t=1$.
Appendix A Additional Discussions on Results
A.1 Additional Qualitative Comparisons
Fig. 11, Fig. 12, Fig. 13, Fig. 14, Fig. 15, Fig. 16, and Fig. 17 provide additional qualitative comparisons of SAR-to-EO image translation results on the QXS-SAROPT [50], SAR2Opt [80], and SpaceNet6 [56] dataset. The GAN-based methods, including Pix2Pix [24] and CycleGAN [84], exhibit severe artifacts due to inherent unstable training in GAN frameworks. Although CFCA-SET [31] and StegoGAN [69] mitigate some of these artifacts, they still produce visually inconsistent results, often failing to preserve fine-grained structural details. Among the LDM-based approaches, BBDM [33] and cBBDM [28] generate smoother outputs and struggle with oversimplified textures and lack of structural alignment due to their direct diffusion process. DGDM [71], on the other hand, relies on an initial translator network to generate SAR-to-EO latents; however, the simplicity of this translator leads to poorly initialized latents, resulting in entirely unrealistic outputs. In contrast, our proposed C-DiffSET effectively addresses these limitations, producing visually coherent and structurally accurate EO images that are closely aligned with the target EO images.
A.2 Progressive Denoising Visualization
Fig. 8 illustrates the progression of generated images across different inference timesteps during the reverse denoising process. Out of the total $T_{\text{test}}=50$ inference steps, we select 11 representative timesteps to visualize the progressive refinement of the output. Using the DDIM [58] noise scheduler, denoising is performed over 50 steps, with timesteps chosen from the range $[1,T]$ to balance sampling efficiency and performance. At earlier timesteps (e.g., $t=981$), the images exhibit significant noise, reflecting the initial latent representation. As the process progresses to later timesteps (e.g., $t=1$), the generated images become increasingly coherent, closely resembling the target EO images.
A.3 Analysis of VAE Reconstruction and Upper Bound Performance
In Fig. 4, we evaluate the reconstruction quality of the VAE encoder $\mathcal{E}_{\text{vae}}$ and decoder $\mathcal{D}_{\text{vae}}$ from the LDM by visualizing their outputs for target EO images and SAR images with varying levels of speckle noise. Our proposed C-DiffSET framework, built upon the LDM architecture, predicts target EO features $\mathbf{z}_{y}=\mathcal{E}_{\text{vae}}(\mathbf{Y})$ embedded in the latent space through the VAE encoder. Thus, the VAE-decoded reconstruction of the target EO feature, $\mathcal{D}_{\text{vae}}(\mathcal{E}_{\text{vae}}(\mathbf{Y}))$, serves as an upper bound for our C-DiffSET, representing the maximum achievable quality given the pretrained LDM. Table 4 quantitatively compares the performance of C-DiffSET with this upper bound. While our framework delivers strong results, especially in terms of perceptual metrics such as LPIPS and FID, the gap to the upper bound highlights the inherent limitations in SAR-to-EO image translation due to noise, misalignments, and the complexities of SAR data. This analysis underscores the potential for future work to narrow this gap, advancing SAR-to-EO image translation toward the limits defined by the upper bound.
  Methods First 50 samples from the SpaceNet6 Dataset
[…]
C-DiffSET (Ours) 85.77 0.132 0.0210 0.546 21.498
$\mathcal{D}_{\text{vae}}(\mathcal{E}_{\text{vae}}(\mathbf{Y}))$ 43.57 0.049 0.1631 0.783 28.633
Table 4: Comparison of C-DiffSET performance with the VAE upper bound on the SpaceNet6 dataset. The upper bound represents the maximum achievable quality defined by the VAE-decoded target EO features.
Figure 9: Impact of total inference steps on performance metrics (FID, LPIPS, SCC, SSIM, and PSNR) and inference time.
Figure 10: Visualization of C-DiffSET results across varying numbers of total inference steps.
Appendix B Further Ablation Studies
B.1 Effect of Total Inference Steps
In Fig. 9 and Fig. 10, we justify the selection of $T_{\text{test}}=50$ inference steps in our experiments. While setting the total inference steps to match the training steps ($T=1,000$) ensures high performance, it incurs prohibitive inference times. To balance performance and efficiency, we evaluate the tradeoff between inference time and metric performance by varying the DDIM noise scheduler’s inference steps. For computational feasibility, experiments are conducted on the first 50 samples of the SpaceNet6 dataset. The results indicate significant performance gains for lower step counts ($T_{\text{test}}=1$ to $T_{\text{test}}=50$), but diminishing returns beyond 50 steps, with negligible improvements in metrics such as FID, LPIPS, and PSNR. Based on this analysis, we select total 50 inference steps as the optimal tradeoff, achieving high-quality outputs within practical inference times (3.376 second per 512$\times$512 image).
Figure 11: Qualitative SAR-to-EO image translation on the QXS-SAROPT dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
Figure 12: Qualitative SAR-to-EO image translation on the QXS-SAROPT dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
Figure 13: Qualitative SAR-to-EO image translation on the QXS-SAROPT dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
Figure 14: Qualitative SAR-to-EO image translation on the SAR2Opt dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
Figure 15: Qualitative SAR-to-EO image translation on the SAR2Opt dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
Figure 16: Qualitative SAR-to-EO image translation on the SpaceNet6 dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
Figure 17: Qualitative SAR-to-EO image translation on the SpaceNet6 dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).